OpenStudy (tanabugg):

Write an equation with a variable on both sides of the equal sign that has infinitely many solutions. Solve the equation and explain why it has an infinite number of solutions.

pooja195 (pooja195):

Both sides of the equations must be the same in order for an equation to have infinite solutions

OpenStudy (tanabugg):

Could x=x be one answer?

pooja195 (pooja195):

No

OpenStudy (anonymous):

that is what i would say, yes

OpenStudy (anonymous):

probably will annoy your teacher but that is the simplest one

pooja195 (pooja195):

:/ I was thinking something along the lines of 8x+4=8x+4

OpenStudy (anonymous):

or to make it fancy you could do what @pooja195 said but make it \[8x+4=4(2x+1)\]

pooja195 (pooja195):

^ right ^-^

OpenStudy (tanabugg):

Thanks a lot guys! I wish I could give you both medals !

pooja195 (pooja195):

Give it to sat :P

OpenStudy (tanabugg):

How could i put that into three sentences?

pooja195 (pooja195):

Both sides of the equations must be the same in order for an equation to have infinite solutions

OpenStudy (anonymous):

i have plenty, pooja deserves one

pooja195 (pooja195):

xD

pooja195 (pooja195):

@tanabugg i did it for u in one :P

OpenStudy (tanabugg):

We are supposed to have three sentences but i think it will work thanks!!

pooja195 (pooja195):

yw :)

OpenStudy (tanabugg):

@pooja195 My teacher graded my work for this problem and said that i need to show my work for the equation and i'm not sure what she means

pooja195 (pooja195):

yea you need to solve for x